### v1

# TRENDY: Gene Regulatory Network Inference Enhanced by Transformer

Abstract
Gene regulatory networks (GRNs) play a crucial role in the control of cellular functions. Numerous methods have been developed to infer GRNs from gene expression data, including mechanism-based approaches, information-based approaches, and more recent deep learning techniques, the last of which often overlooks the underlying gene expression mechanisms. In this work, we introduce TRENDY, a novel GRN inference method that integrates transformer models to enhance the mechanism-based WENDY approach. Through testing on both simulated and experimental datasets, TRENDY demonstrates superior performance compared to existing methods. Furthermore, we apply this transformer-based approach to three additional inference methods, showcasing its broad potential to enhance GRN inference.
1 Introduction
The expression of genes can be regulated by other genes. We can construct a directed graph, where vertices are genes, and edges are regulatory relationships. This graph is called a gene regulatory network (GRN). GRNs are essential to understanding the mechanisms governing cellular processes, including how cells respond to various stimuli, differentiate, and maintain homeostasis [1, 2]. Understanding the GRN structure is important for developmental biology [3, 4, 5, 6], disease research [7, 8, 9], and even studying macroscopic behavior [10, 11, 12].
It is difficult to directly determine the GRN structure with experiments. Instead, researchers develop methods that infer the GRN structure through gene expression data. Some methods, such as WENDY [13] and NonlinearODEs [14], are mechanism-based: they construct dynamical models for gene expression, and fit with expression data to determine the GRN [15, 16]. Some methods, like GENIE3 [17] and SINCERITIES [18], are information-based: they directly find genes that contain information (predictability) for the target gene, and assume that such information implies regulation [19, 20].
Recently, there are some deep learning-based methods that infer GRN with different neural network frameworks, such as convolutional neural networks [21], recurrent neural networks [22], variational autoencoders [23], graph neural networks [24, 25], etc. Deep learning frameworks can perform well on many tasks. However, because of the large number of tunable parameters in a neural network, it is difficult to explain why it works. These deep learning-baed GRN inference methods have similar problems: they generally apply popular neural networks as black boxes without integration with biological mechanisms, and thus lacking interpretability.
The transformer model is a deep learning architecture designed to handle sequential data [26]. It can capture long-range dependencies and relationships through the self-attention mechanism. The encoder layer of the transformer model can process the input through a neural network, so that the output is close to the given target.
Some researchers have developed GRN inference methods based on the transformer model [27] or its variants [28]. The STGRNS method [27] predicts the existence of regulation between two genes solely based on the expression levels of these two genes. Since other genes are not considered, this approach makes it difficult to distinguish between direct regulation ($G_{i}\to G_{j}$) and indirect regulation through a third gene ($G_{i}\to G_{k}\to G_{j}$). The GRN-transformer method [28] uses multiple inputs, including the GRN inferred by an information-based method, PIDC [29], but it also does not consider the biological dynamics of gene regulation.
To combine the powerful deep learning techniques and biological understanding of gene regulation, we propose a novel GRN inference method that applies transformer models to enhance a mechanism-based GRN inference method, WENDY. This new method, TRansformer-Enhanced weNDY, is abbreviated as TRENDY. WENDY method is based on a dynamical model for the covariance matrix of genes. In TRENDY, we first use a transformer model to construct a pseudo-covariance matrix that performs better in WENDY. Then we apply another transformer model that directly enhances the inferred GRN.
The idea for the second half of TRENDY can be used to enhance the inferred results by other GRN inference methods. We apply transformer models to three GRN inference methods, GENIE3, SINCERITIES, and NonlinearODEs, to obtain their enhanced versions, GENIE3-rev, SINCERITIES-rev, and NonlinearODEs-rev.
We test four transformer-enhanced methods against their non-transformer counterparts on two simulated data sets and two experimental data sets. All four transformer-enhanced methods outperform existing methods, and TRENDY ranks the first among all methods.
In order to train a deep learning model, such as the transformer model used in this paper, we need a sufficient amount of training data. Specifically, since we want the GRN inference method to work well in different situations, the training data should contain many different data sets that correspond to different GRNs. In reality, there are not many available experimental gene expression data sets with known GRNs. Therefore, we consider generating synthetic data sets.
To simulate new gene expression data, the most common approach is to assume that gene expression follows a certain mechanism, such as an ordinary differential equation (ODE) system [30, 31] or a stochastic differential equation (SDE) system [32, 33]. After determining the parameters of this system, one can simulate a discretized version of this system and obtain new data. Readers may refer to a review [34] for this mechanism approach. Another approach is to apply deep learning to learn from experimental gene expression data and generate new data that mimic the existing data. Such generative neural networks can be variational autoencoders [23] or generative adversarial networks [35].
The advantage of the differential equation approach is that one can change the parameters to generate new data that correspond to another GRN, while the deep learning approach can only reproduce existing data. The disadvantage of the differential equation approach is that the modeled mechanism might not be a perfect fit to reality. Therefore, the generated data might be questionable, especially those generated by an oversimplified model, such as a linear ODE system. Instead, the deep learning approach can generate new data that are indistinguishable from the experimental data.
In this paper, we need to feed the transformer models with data from many different GRNs. Therefore, we adopt the mechanism approach with a frequently used nonlinear and stochastic system [36, 18, 13]:
$\mathrm{d}X_{j}(t)=V\left\{\beta\prod_{i=1}^{n}\left[1+(A_{\mathrm{true}})_{i,j}\frac{X_{i}(t)}{X_{i}(t)+1}\right]-\theta X_{j}(t)\right\}\mathrm{d}t+\sigma X_{j}(t)\mathrm{d}W_{j}(t).$ (1)
The matrix $A_{\mathrm{true}}$ is a randomly generated ground truth GRN, where $(A_{\mathrm{true}})_{i,j}>0/=0/<0$ means that gene $i$ has positive/no/negative regulatory effects on gene $j$. $X_{i}(t)$ is the expression level of gene $i$ at time $t$, and $W_{j}(t)$ is a standard Brownian motion, independent with other $W_{i}(\tau)$. The parameter values are $V=30$, $\beta=1$, $\theta=0.2$, and $\sigma=0.1$. Our goal is to infer $A_{\mathrm{true}}$ from $X_{j}(t)$.
Most deep learning-based GRN inference methods are trained on experimental data sets, whose amount is very limited. Therefore, such methods need to carefully choose the neural network structure and the training procedure under this limitation. In this paper, we consider a scenario in which the training data set is sufficiently large, and study how deep learning can achieve the best performance without data limitation.
2 Results
2.1 TRENDY method
WENDY method [13] uses single-cell gene expression data measured at two time points, each of size $m\times n$ (mRNA counts of $n$ genes for $m$ cells). Here we do not know how cells at different time points correspond to each other, since the measurement of single-cell level gene expression needs to kill the cells [37]. For data at two time points, the cell numbers $m$ can be different.
For gene expression data at time points 00 and $t$, WENDY applies graphical lasso [38] to calculate the covariance matrices for different genes, $K_{0}$ and $K_{t}$, each of size $n\times n$. We can construct a general SDE system like Eq. 1 to model the dynamics of gene regulation. After linearization of this system, one can obtain an approximated relation for $K_{0},K_{t}$, and the ground truth GRN $A_{\mathrm{true}}$, also of size $n\times n$:
$K_{t}=(I+tA_{\mathrm{true}}^{\mathrm{T}})K_{0}(I+tA_{\mathrm{true}})+D+E.$ (2)
Here $I$ is an $n\times n$ identity matrix, $D$ is an unknown $n\times n$ diagonal matrix, and $E$ is the unknown error introduced by linearization. Then WENDY solves the GRN matrix $A$ as a non-convex optimization problem:
$\arg\min_{A}f(A):=\frac{1}{2}\sum_{i\neq j}\{[K_{t}-(I+tA^{\mathrm{T}})K_{0}(I+tA)]_{i,j}\}^{2}.$ (3)
Due to the unknown diagonal matrix $D$, this summation does not count elements with $i=j$. The result of this optimization problem is denoted as
$A_{0}=\text{WENDY}(K_{0},K_{t}).$
If WENDY works perfectly, then
$K_{t}=(I+tA_{0}^{\mathrm{T}})K_{0}(I+tA_{0})+D,$
meaning that $(I+tA_{\mathrm{true}}^{\mathrm{T}})K_{0}(I+tA_{\mathrm{true}})$ and $(I+tA_{0}^{\mathrm{T}})K_{0}(I+tA_{0})$ differ by $E$, the error introduced by the linearization, and $A_{0}$ may not match $A_{\mathrm{true}}$ accurately. If we want to derive a more accurate and complicated equation of $K_{0}$, $K_{t}$, and $A_{\mathrm{true}}$ than Eq. 2, this equation might be numerically unstable to solve. Therefore, we need to find another way to improve the WENDY method.
Define
$K_{t}^{*}=(I+tA_{\mathrm{true}}^{\mathrm{T}})K_{0}(I+tA_{\mathrm{true}}).$ (4)
If we replace $K_{t}$ by $K_{t}^{*}$ in Eq. 3 and solve $A$, then this inferred $A$ should be very close to $A_{\mathrm{true}}$. However, in practice, we only have $K_{0}$ and $K_{t}$, but not $K_{t}^{*}$, since $A_{\mathrm{true}}$ is unknown.
The core idea is to train a transformer model $\mathrm{TE}(k=1)$ with input $K_{t}$, so that the output $K_{t}^{\prime}$ is close to $K_{t}^{*}$. Then we can calculate
$A_{1}=\text{WENDY}(K_{0},K_{t}^{\prime}),$
which is closer to $A_{\mathrm{true}}$ than $A_{0}$. We can further improve $A_{1}$ using another transformer model $\mathrm{TE}(k=3)$, along with the information in $K_{0},K_{t}$, so that the output $A_{2}$ is even closer to $A_{\mathrm{true}}$. These two transformer models will be discussed in the next subsection.
Fig. 1 describes the training and testing workflow for TRENDY. To train TRENDY, one needs to
(1) generate many random $A_{\mathrm{true}}$ and their corresponding gene expression data;
(2) calculate the covariance matrices $K_{0}$ and $K_{t}$, and then calculate $K_{t}^{*}$ from $K_{0},K_{t},A_{\mathrm{true}}$;
(3) use $K_{t}$ as input and $K_{t}^{*}$ as target to train the $\mathrm{TE}(k=1)$ model, and then use the trained $\mathrm{TE}(k=1)$ model to calculate $K_{t}^{\prime}$ from $K_{t}$;
(4) use WENDY to calculate $A_{1}$ from $K_{0},K_{t}^{\prime}$;
(5) use $A_{1},K_{0},K_{t}$ as input and $A_{\mathrm{true}}$ as target to train the $\mathrm{TE}(k=3)$ model.
For applying TRENDY to infer GRN from existing data, one needs to omit step (1) and
(2) calculate the covariance matrices $K_{0}$ and $K_{t}$;
(3) use the trained $\mathrm{TE}(k=1)$ model to calculate $K_{t}^{\prime}$ from $K_{t}$;
(4) use WENDY to calculate $A_{1}$ from $K_{0},K_{t}^{\prime}$;
(5) feed $A_{1},K_{0},K_{t}$ to the trained $\mathrm{TE}(k=3)$ model to calculate the final inferred GRN $A_{2}$, the final output of TRENDY.
See https://github.com/YueWangMathbio/TRENDY for all code files and data used in this paper.
data
generator
expression
data
expression
data
[…]
revised
covariance
inferred
GRN
inferred
GRN
[…]
covariance
matrix
covariance
matrix
$\mathrm{TE}(k=1)$
model
$\mathrm{TE}(k=3)$
model
WENDY
method
Figure 1: Workflow of training and testing the TRENDY method. Rectangles with grids are numerical matrices, where the color scale represents the numerical value. Circles are mechanisms that take in matrices and produce matrices. Single and double arrows represent inputs and outputs of each mechanism, and dashed arrows represent direct calculations. To apply TRENDY after training, the step in the dashed rectangle is omitted.
2.2 Transformer model $\mathrm{TE}(k)$
We build a deep learning framework $\mathrm{TE}(k)$ that is based on transformer encoder layers, where $k$ is the major hyperparameter that describes the number of input matrices. Besides, there are three hyperparameters that can be tuned: the model dimension $d$; the number of encoder layers $l$; the number of attention heads $h$ of the encoder layer. We use this $\mathrm{TE}(k)$ model with different hyperparameters in TRENDY and other transformer-enhanced methods. See Table 1 for the general structure of $\mathrm{TE}(k)$, along with the shape of data after each layer. Notice that the same transformer encoder layer can handle inputs of different lengths. This means that the gene number $n$ is not a predetermined hyperparameter, and we do not need to train a different model for each $n$. We will train the $\mathrm{TE}(k)$ model with $n=10$ genes and test on data sets with $n=10/18/20$ genes.
For the first half of the TRENDY method, the $\mathrm{TE}(k=1)$ model has $k=1$ input matrix $K_{t}$, $d=64$ model dimension, $l=7$ encoder layers, and $h=4$ attention heads. For the second half of the TRENDY method, the $\mathrm{TE}(k=3)$ model has $k=3$ input matrices of the same size, $A_{1},K_{0},K_{t}$, $d=64$ model dimension, $l=7$ encoder layers, and $h=8$ attention heads.
input ($k$ matrices) $k$ groups of $n\times n$
linear embedding layer ($1$ to $d$) $k$ groups of $n\times n\times d$
segment embedding layer (omitted when $k=1$) $k$ groups of $n\times n\times d$
2-D positional encoding layer $k$ groups of $n\times n\times d$
flattening and concatenation $(n^{2})\times(dk)$
$l$ layers of transformer encoder with head number $h$ $(n^{2})\times(dk)$
linear embedding layer ($(dk)$ to $1$) $n^{2}\times 1$
reshaping and output $n\times n$
training target $n\times n$
Table 1: Structure of the $\mathrm{TE}(k)$ model. The left column lists all layers. The right column illustrates the shape of data after each layer, where $n$ is the number of genes, $d$ is the model dimension, and $k$ is the number of input matrices. Notice that $k,d,l,h$ are tunable hyperparameters, but $n$ is not a predetermined hyperparameter.
2.3 Other transformer-enhanced methods
GENIE3 [17] works on single-cell expression data at one time point. It uses random forest to select genes that have predictability for the target gene, thus being information-based.
SINCERITIES [18] works on single-cell expression data at multiple time points. It runs regression on the difference between distributions of expression levels, also being information-based.
NonlinearODEs [14] works on bulk expression data at multiple time points. This method fits the data with a nonlinear differential equation system, which is thus mechanism-based.
For each of these three methods, we train a $\mathrm{TE}(k)$ model to enhance the inferred results, similar to the $\mathrm{TE}(k=3)$ model in the second half of TRENDY. Here each $\mathrm{TE}(k)$ model has $d=64,l=7,h=8$, and the training target is $A_{\mathrm{true}}$.
For inferred GRN $A_{\mathrm{G}}$ by GENIE3, we train a $\mathrm{TE}(k=2)$ model with $A_{\mathrm{G}}$ and the covariance matrix $K$ as input. Then we can use this trained model to calculate a more accurate $A_{\mathrm{G}}^{\prime}$ from $A_{\mathrm{G}}$ and $K$. This method is named GENIE3-rev.
For inferred GRN $A_{\mathrm{S}}$ by SINCERITIES, we train a $\mathrm{TE}(k=1)$ model with $A_{\mathrm{S}}$ as input, and name it SINCERITIES-rev.
For inferred GRN $A_{\mathrm{N}}$ by NonlinearODEs, we train a $\mathrm{TE}(k=1)$ model with $A_{\mathrm{N}}$ as input, and name it NonlinearODEs-rev.
2.4 Performance of different methods
The same as the WENDY paper [13], we test the performance of different GRN inference methods on two simulated data sets and two experimental data sets. To compare the inferred GRN $A_{\mathrm{pred}}$ and the ground truth GRN $A_{\mathrm{true}}$, we adopt the common practice that calculates two area under curve (AUC) scores: AUROC and AUPRC [13]. Both AUC scores are between 00 and $1$, where $1$ means perfect matching, and 00 means perfect mismatching.
2.4.1 SINC data sets
We use the same method for generating training data to generate 1000 testing samples, and name this data set as SINC [18, 13]. Each sample with a random $A_{\mathrm{true}}$ has 100 cells measured at time points $0.0,0.1,\ldots,1.0$.
For WENDY and TRENDY, we set the first time point to $0.0$ and vary the second time point from $0.1$ to $1.0$. For GENIE3 and GENIE3-rev, we set the time point to any of $0.1,\ldots,1.0$. For SINCERITIES, SINCERITIES-rev, NonlinearODEs, and NonlinearODEs-rev, we use all time points. See Fig. 2 and Table 2 for the results.
Figure 2: Average AUROC and AUPRC values of eight GRN inference methods on the SINC data set. Notice that SINCERITIES, NonlinearODEs and their revised versions need data at all time points.
2.4.2 DREAM4 data set
DREAM4 data set is also generated by simulating an SDE system, and consists of multiple data sets under different settings [39]. DREAM4 is commonly used as a benchmark for GRN inference [13, 18]. Here we use the 10-gene time series data set. There are five $A_{\mathrm{true}}$, each corresponding to the levels of 10 genes for five cells at 21 time points.
For WENDY and TRENDY, we consider any pair of time points and take average. For GENIE3 and GENIE3-rev, we consider any one time point and take average. For SINCERITIES, SINCERITIES-rev, NonlinearODEs, and NonlinearODEs-rev, we use any consecutive 11 time points and take average. See Fig. 3 and Table 2 for the results.
Figure 3: Average AUROC and AUPRC values of eight GRN inference methods on the DREAM4 data set.
2.4.3 THP-1 data set
THP-1 data set measures monocytic THP-1 human myeloid leukemia cells at eight time points [40]. Each time point has 120 cells measured. The same as previous papers [18, 13], we consider the expression levels of 20 genes.
For WENDY and TRENDY, we consider any pair of time points and take average. For GENIE3 and GENIE3-rev, we consider any one time point and take average. For SINCERITIES, SINCERITIES-rev, NonlinearODEs, and NonlinearODEs-rev, we use all eight time points. See Fig. 4 and Table 2 for the results.
Figure 4: AUROC and AUPRC values of eight GRN inference methods on the THP-1 data set.
2.4.4 hESC data set
hESC data set measures human embryonic stem cell-derived progenitor cells at six time points [41]. Each time point has 66–172 cells measured. The same as in a previous paper [13], we consider the expression levels of 18 genes.
For WENDY and TRENDY, we consider any pair of time points and take average. For GENIE3 and GENIE3-rev, we consider any one time point and take average. For SINCERITIES, SINCERITIES-rev, NonlinearODEs, and NonlinearODEs-rev, we use all six time points. See Fig. 5 and Table 2 for the results. Notice that the ground truth GRN $A_{\mathrm{true}}$ has very few nonzero values. This is why the AUPRC value is much lower than the AUROC value.
Figure 5: AUROC and AUPRC values of eight GRN inference methods on the hESC data set.
2.4.5 Overall evaluation
We summarize the performance of all eight methods on four data sets. See Table 2 for details. The four transformer-enhanced methods (TRENDY, GENIE3-rev, SINCERITIES-rev, NonlinearODEs-rev) are all better than the four known methods. This means that our idea of enhancing GRN inference methods with $\mathrm{TE}(k)$ models works well universally. Although the training data only has $n=10$ genes, and the THP-1 and hESC data sets have $n=20$ and $n=18$ genes, the overall performance of transformer-enhanced methods is better than their non-transformer counterparts.
The core method of this paper, TRENDY, ranks the first among all eight methods. Besides, the TRENDY method is better than the WENDY method on three data set, and has almost the same performance as WENDY on one data set (THP-1). The other three transformer-enhanced methods are worse than their non-transformer counterparts on at least one data set. This means that TRENDY’s performance is robust on different data sets. In sum, TRENDY has satisfactory performance among all methods tested.
WEN TRE GEN GENr SIN SINr NON NONr
S-RO 0.6654 0.8547 0.4062 0.8703 0.6783 0.7964 0.5076 0.5976
S-PR 0.6109 0.7519 0.5061 0.7672 0.5829 0.6637 0.5313 0.5658
D-RO 0.4899 0.5341 0.5636 0.4589 0.4908 0.4995 0.4806 0.5712
D-PR 0.2080 0.2177 0.2286 0.1799 0.1919 0.2034 0.1705 0.2452
T-RO 0.5261 0.5557 0.4484 0.5506 0.6261 0.5251 0.5338 0.4808
T-PR 0.3972 0.3669 0.3546 0.3781 0.3852 0.3412 0.3486 0.3302
h-RO 0.4997 0.5311 0.5913 0.6008 0.4198 0.4871 0.5971 0.6233
h-PR 0.0392 0.0376 0.0468 0.0435 0.0261 0.0294 0.0534 0.0641
Total 3.4364 3.8497 3.1456 3.8493 3.4011 3.5458 3.2229 3.4782
Rank 5th 1st 8th 2nd 6th 3rd 7th 4th
Table 2: Performance summary of eight GRN inference methods on four data sets. For row names, S, D, T, h stand for SINC, DREAM4, THP-1, hESC data sets; -RO and -PR represent AUROC and AUPRC values. For column names, WEN, TRE, GEN, GENr, SIN, SINr, NON, NONr represent WENDY, TRENDY, GENIE3, GENIE3-rev, SINCERITIES, SINCERITIES-rev, NonlinearODEs, NonlinearODEs-rev. The “Total” row is the sum of all eight AUC values, and the last row is the rank of each method by the total score. Red numbers indicate that they are smaller than their non-transformer counterparts (left neighbors) by at least 0.01, and blue color means larger by at least 0.01. For instance, TRENDY has S-RO$=0.8547$, larger than WENDY’s S-RO$=0.6654$ by at least 0.01, so that $0.8547$ is in blue.
3 Discussion
In this paper, we present the TRENDY method, which uses the transformer model $\mathrm{TE}(k)$ to enhance the mechanism-based GRN inference method WENDY. TRENDY is tested against three other transformer-enhanced methods and their original forms, and ranks the first. This work explores the potential of deep learning methods in GRN inference when there are sufficient training data. Besides, TRENDY is developed on the biological dynamics of gene regulation, and it is more interpretable than other deep learning-based methods that treat GRN inference as a pure prediction task.
The essential difficulty of GRN inference is the lack of data with experimentally verified GRNs. Many deep learning models need significantly more data with known GRNs, and we have to simulate new data from a generator. Since we do not fully understand the relation between the dynamics of gene expression and the ground truth GRN, it is difficult to evaluate the quality of generated data along with a random GRN. We also cannot guarantee that GRN inference methods that perform well on known experimental data sets can still be advantageous if there are many more experimental data with different GRNs.
We train the models with data of $n=10$ genes, and they work well on data with $n=10/18/20$ genes. If we want the models to work on data with many more genes, we need to train the models with corresponding data. However, when the gene number $n$ is large, the number of different GRNs increases exponentially, and the number of training samples should also increase violently. The time cost for generating such data can be quite long. Even with the same amount of training samples, the training speed of the $\mathrm{TE}(k)$ model is proportional to $n^{4}$, since the input length of the transformer encoder layer is proportional to $n^{2}$, and it needs to calculate the attention scores between all pairs of input tokens. Therefore, scaling TRENDY for a much larger $n$ is extremely time-consuming.
In this paper, the training data are generated by Eq. 1, which only describes the dynamics of mRNA count as a continuous-state process. This might be an oversimplification of the gene regulation in reality, where the gene state and the protein count also should be considered [42]. Besides, the experimental gene expression data often suffer from incomplete measurement, where many mRNAs are not recorded, and there are many missing values. Therefore, when we train on perfectly measured data and test on data with missing values, the performance is not guaranteed. We can manually add measurement errors and missing values to the training data to solve this problem.
In our testing, we find that using $\mathrm{TE}(k=1)$ model to directly enhance an inferred GRN, like SINCERITIES-rev or NonlinearODEs-rev, may not be sufficiently stable. Instead, using $\mathrm{TE}(k>1)$ model with both inferred GRN and corresponding covariance matrices, like TRENDY or GENIE3-rev, can achieve better results, since the covariance matrices may contain extra information besides that contained in the inferred GRN.
Some deep learning-based GRN inference methods directly use the raw gene expression data as the model input. We think that the covariance matrix contains most of the information regarding the relation between any two genes, and the covariance matrix is calculated from graphical lasso, thus being more regular than the original expression data. Therefore, using covariance matrices in deep learning GRN inference methods should be a promising approach.
4 Methods
4.1 Training data
To generate training data for our TRENDY method, we use Eq. 1 to numerically simulate gene expression data with the Euler-Maruyama method [43] for $n=10$ genes. Every time we generate a random ground truth GRN $A_{\mathrm{true}}$, where each element has probability $0.1/0.8/0.1$ to be $-1/0/1$. Then we use Eq. 1 with this random $A_{\mathrm{true}}$ to run the simulation from time $0.0$ to time $1.0$ with time step $0.01$ to obtain one trajectory, where the initial state at time 00 is random. This is repeated 100 times to obtain 100 trajectories, similar to the situation where we measure the single-cell expression levels of $n$ genes for 100 cells [44]. We only record expression levels at time points $0.0,0.1,0.2,\ldots,1.0$, and then calculate the covariance matrix for $n=10$ genes at each time point with graphical lasso.
We repeat the simulation to obtain $1.01\times 10^{5}$ samples, where $10^{5}$ samples are for training, and $10^{3}$ samples are for validation (used in hyperparameter tuning and early stopping). Since the TRENDY method only needs two time points, we fix the first time point to be $0.0$, and the second time point can be $0.1,0.2,\ldots,1.0$. Therefore, each simulated sample corresponds to 10 pairs of $K_{0}$ and $K_{t}$. This means that the training set has $10^{6}$ pairs of $K_{0}$ and $K_{t}$, where each $t$ in $0.1,0.2,\ldots,1.0$ corresponds to $10^{5}$ pairs.
For GENIE3-rev, we consider each time point $t$ in $0.1,0.2,\ldots,1.0$ and calculate the inferred GRN $A_{\mathrm{G}}$. Then there are $10^{6}$ pairs of $A_{\mathrm{G}}$ and $K_{t}$ for training the $\mathrm{TE}(k=2)$ model. For SINCERITIES-rev and NonlinearODEs-rev, we use the data at all time points $0.0,0.1,0.2,\ldots,1.0$ to infer the GRN $A_{\mathrm{S}}$ and $A_{\mathrm{N}}$. Therefore, each method has $10^{5}$ samples for training the $\mathrm{TE}(k=1)$ model.
4.2 Loss function
We use AUROC and AUPRC to measure the difference between $A_{\mathrm{true}}$ and the inferred $A_{\mathrm{pred}}$. Since $A_{\mathrm{true}}$ can take three values $-1/0/1$, the calculation of AUROC and AUPRC need some revisions [13]. Unfortunately, for fixed $A_{\mathrm{true}}$ and arbitrary $A_{\mathrm{pred}}$, AUROC and AUPRC can only take finitely many values, and are not continuous functions of $A_{\mathrm{pred}}$. We cannot train a neural network with a non-continuous loss function, since it is not differentiable, and the back propagation cannot work. Therefore, we cannot directly use AUROC or AUPRC as the loss function when training $\mathrm{TE}(k)$ models. There have been some surrogate loss functions for AUROC or AUPRC [45, 46], but we find that with enough training data, the mean square error (MSE) function can lead to satisfactory training results. We use MSE as the loss function for the $\mathrm{TE}(k=1)$ model in TRENDY. For all other $\mathrm{TE}(k)$ models used, the loss function is also MSE, but the diagonal elements are ignored, since they represent the autoregulation of one gene $i$ to itself, which cannot be inferred by the methods discussed in this paper [47].
4.3 Segment embedding
For $\mathrm{TE}(k)$ model with $k>1$ input matrices, after processing them with the linear embedding layer, we need to add segment embeddings to them. We give the segment ID $0/1/2/\ldots$ to the first/second/third/$\ldots$ group of processed inputs, and use an embedding layer to map each segment ID to a $d$-dimensional embedding vector. Repeat this embedding vector $n\times n$ times to obtain an $(n\times n\times d)$ array, and add it to the $(n\times n\times d)$ arrays after the linear embedding layer.
4.4 Positional encoding
For the $\mathrm{TE}(k)$ model, the 2-dimensional positional encoding layer [27] adds an $(n\times n\times d)$ array PE to each of the embedded input. For $x$ and $y$ in $1,2,\ldots,n$ and $j$ in $41,2,\ldots,d/$ 4, PE is defined as
$\begin{split}&\text{PE}[x,y,2j-1]=\text{PE}[x,y,2j-1+d/2]=\sin[(y-1)\times 10^{-32(j-1)/d}],\\ &\text{PE}[x,y,2j]=\text{PE}[x,y,2j+d/2]=\cos[(y-1)\times 10^{-32(j-1)/d}],\\ &\text{PE}[x,y,2j-1+d/4]=\text{PE}[x,y,2j-1+3d/4]=\sin[(x-1)\times 10^{-32(j-1)/d}],\\ &\text{PE}[x,y,2j+d/4]=\text{PE}[x,y,2j+3d/4]=\cos[(x-1)\times 10^{-32(j-1)/d}].\end{split}$ (5)
In our numerical simulations, we find that this 2-dimensional positional encoding is crucial for the $\mathrm{TE}(k)$ model to produce better results.
4.5 Training setting and cost
All $\mathrm{TE}(k)$ models used in this paper are trained for 100 epochs. For each epoch, we evaluate the model performance on the validation data set. If the performance does not improve for 10 consecutive epochs, we stop training early. For all transformer encoder layers, the dropout rate is 0.1. The optimizer is Adam with learning rate 0.001.
Data generation and model training are conducted on a desktop computer with Intel i7-13700 CPU and NVIDIA RTX 4080 GPU. Data generation takes about one week with CPU parallelization. Training for each $\mathrm{TE}(k)$ model takes about two hours with GPU acceleration. After training, the time cost of applying TRENDY on a data set with $\sim 10$ genes and $\sim 100$ cells is under one second.
Acknowledgments
The authors would like to thank Dr. Zikun Wang for helpful comments. YW would like to thank Dr. Mingda Zhang, Dr. Lingxue Zhu, and Mr. Vincent Zhang for an intriguing discussion. YP would like to thank Irving Institute for Cancer Dynamics of Columbia University for hosting the Summer Research Program.
Data and code availability
Main function of TRENDY method, including a tutorial and the training files, along with other data and code files used in this paper, can be found in
https://github.com/YueWangMathbio/TRENDY